# Condensing Against Online Adversaries

### v1

@@ Abstract
We investigate the task of deterministically condensing randomness from Online Non-Oblivious Symbol Fixing (oNOSF) sources, a natural model of defective random sources for which it is known that extraction is impossible [AORSV, EUROCRYPT’20]. A $(g,\ell)$-oNOSF source is a sequence of $\ell$ blocks $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})\sim(\{0,1\}^{n})^{\ell}$, where at least $g$ of the blocks are good (are independent and have some min-entropy), and the remaining bad blocks are controlled by an online adversary where each bad block can be arbitrarily correlated with any block that appears before it.
The existence of condensers was recently studied in [CGR, FOCS’24]. They proved condensing impossibility results for various values of $g$ and $\ell$, and they showed the existence of condensers matching the impossibility results in the special case when $n$ is extremely large compared to $\ell$ (i.e., the setting of few blocks of large length).
@@
Finally, to understand the case of small $n$, we focus on $n=1$ which corresponds to online non-oblivious bit-fixing (oNOBF) sources. We introduce and initiate a systematic study of a new, natural notion of the influence of Boolean functions, which we call online influence, and believe is of independent interest. Using tools from Boolean Fourier analysis, we establish tight bounds on the total online influence of Boolean functions, which imply extraction lower bounds. Several problems remain open regarding this new measure of influence; progress on these will lead to improved extractors and condensers for oNOBF sources or further strengthen our lower bounds.
@@ 1 Introduction
With this, we are ready to formally define condensers:
@@
When $\mathcal{X}$ is the family of all distributions, it is folklore that no non-trivial condensing is possible. [^1] Assuming $m\leq n$ (wlog this holds since $\left\lvert\mathsf{Cond}(\{0,1\}^{n})\right\rvert\leq 2^{n}$), $m-k_{out}\geq(n-k_{in})-\log(1/(1-\varepsilon))$ and hence the output entropy rate cannot be more than the input entropy rate without incurring extremely large error ($>0.999$). So, we additionally assume that $\mathcal{X}$ is a structured family of sources. [^2] A different route, that has been widely studied, is to assume access to a short independent seed. In this work, we will limit ourselves to the seedless setting. Since extractors are the highest quality condensers, a significant amount of work has focused on constructing extractors for interesting family of sources, such as: sources generated by small circuits, two independent sources, algebraically generated sources, sources generated by small space sources, and many more [tv00, chattopadhyay_explicit_2019, dgw09extractpolynomials, kamp_deterministic_2007].
@@
In this work, we focus on one natural family of sources where it is known that extraction is impossible. The family we consider are known as online non-oblivious symbol fixing sources (oNOSF sources ). [^3] These sources are in contrast to non-oblivious symbol fixing (NOSF) sources where bad blocks can be arbitrary functions of all the good blocks. These sources were introduced in [chor_bit_1985] with applications in leakage-resilient cryptography, and have been well-studied. Formally:
@@
The remainder of our introduction is structured as follows. We give an overview of previous work in Section 1.1 before presenting our main existential and explicit condenser results in Section 1.2. In Section 1.3, we show how our results have implications for collective coin flipping and sampling protocols. Next, we introduce our notion of online influence and related results in Section 1.4. We end by defining a local version of oNOSF sources and give explicit extractors for them in Section 1.5.
@@ 1.1 Previous work
The study of condensers for oNOSF sources was initiated by [aggarwal_how_2020]. [^4] In [aggarwal_how_2020], these sources were called SHELA (Somewhere Honest Entropic Look Ahead) sources. Their results include the following:
  • It is impossible to extract from uniform oNOSF sources (even when the fraction of good blocks is an arbitrary constant).
@@
oNOSF sources were further studied by [CGR_seedless_condensers], where they obtained the following results:
@@
We also mention a related family of sources, namely adversarial Chor-Goldreich sources. uniform oNOSF sources can be seen as a special case of adversarial Chor-Goldreich sources where the good blocks are uniform. Constructing condensers where the output entropy rate is $g/\ell$ for adversarial Chor-Goldreich sources is already a challenging task, although such condensers in various parameter regimes have been recently constructed [doron_almost_2023, GLZ_cg_condenser]. [DMOZ24condenseunpredictability] recently constructed condensers for a related more general model.
@@ 1.2.1 Existential condensers
For all $\ell,\varepsilon$ where $\ell\geq O(\log(1/\varepsilon))$, and $n=10^{4}$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.51\ell,\ell,n\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq 0.99m$ where $m=O(\ell+\log(1/\varepsilon))$. Furthermore, when $n=\omega(1)$, the output entropy rate becomes $1-o(1)$.
@@ 1.2.2 Explicit condensers
For all $n,\varepsilon$ and constant $\ell$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.5\ell+1,\ell\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(m/\varepsilon))$ where $m=\Omega(n)$.
@@ 1.2.3 Transforming low-entropy oNOSF sources to uniform oNOSF sources
For all $\ell,n,k,\varepsilon$ where $n=\operatorname{poly}(\log(\ell)),k=O(\log(\ell/\varepsilon))$, there exists a function $f$ such that $f$ transforms ${\left(0.51\ell,\ell,n,k\right)}$-oNOSF sources into uniform ${\left(0.509\ell,\ell,m\right)}$-oNOSF sources with error $\varepsilon$ where $m=O(k)$.
@@ 1.3 Application to collective coin flipping and collective sampling
Condensing from oNOSF sources can be viewed as a special case of coin flipping and collective sampling protocols in the full information model that arise in fault-tolerant distributed computing.
@@ 1.3.1 Background
One way to allow for more corruptions (almost linear) amongst players is to consider “protocols” that allow more rounds of communication. In particular, a protocol can be thought of as a tree where each vertex represents a “round” where in every round the following happens: all good players sends their bits, then all bad players send their bits as a function of the bits of the good players, and they jointly compute a function of these bits. Depending on the outcome of the function, everyone branches on one branch in this tree. Furthermore, every leaf is labelled with final outcomes (say 00 or $1$) and, once you reach a leaf, that is the outcome that everybody agrees on. [ggl98] initiated the study of protocols where the outcomes are from a larger range and where the bad players are trying to minimize the largest probability of any outcome. They called this problem “collective sampling.”
@@ 1.3.2 Known results
[ben-or_collective_1989] showed that for protocols with outcomes $\{0,1\}$, $b$ bad players can always ensure some outcome occurs with probability at least $\frac{1}{2}+\frac{b}{2\ell}$. [alonnoar93collective] first constructed a protocol that can handle $O(n)$ corruptions. Follow-up works tried to reduce the number of rounds in this protocol where, in some settings, players were allowed to send more bits per round [russellzuckerman01, feige_noncryptographic_1999].
[ggl98] showed that for all collective sampling protocols and all outcomes, there exists a way for $b$ bad players to coordinate and ensure that an outcome that happens without corruption with probability $p$, now happens with probability $p^{1-(b/n)}\geq p\left(1+\frac{b}{n}\log(1/p)\right)$. For further results and bounds, see [dodis2006fault].

### v2

# Condensing and Extracting Against Online Adversaries

@@ Abstract
We investigate the tasks of deterministically condensing and extracting randomness from Online Non-Oblivious Symbol Fixing (oNOSF) sources, a natural model of defective random sources for which it is known that extraction is impossible in many parameter regimes [AORSV, EUROCRYPT’20]. A $(g,\ell)$-oNOSF source is a sequence of $\ell$ blocks $\mathbf{X}=(\mathbf{X}_{1},\dots,\mathbf{X}_{\ell})\sim(\{0,1\}^{n})^{\ell}$, where at least $g$ of the blocks are good (are independent and have some min-entropy), and the remaining bad blocks are controlled by an online adversary where each bad block can be arbitrarily correlated with any block that appears before it.
The existence of condensers (in regimes where extraction is impossible) was recently studied in [CGR, FOCS’24]. They proved condensing impossibility results for various values of $g$ and $\ell$, and they showed the existence of condensers matching the impossibility results in the special case when $n$ is extremely large compared to $\ell$ (i.e., the setting of few blocks of large length).
@@
Next, we turn to understanding the possibility of extraction from oNOSF sources. For proving lower bounds, we introduce and initiate a systematic study of a new, natural notion of the influence of functions, which we call online influence, and believe is of independent interest. Using tools from Fourier analysis, we establish tight bounds on the total online influence of functions, which imply extraction lower bounds. Lastly, we give explicit extractor constructions for oNOSF sources, using novel connections to leader election protocols, and further constructing the required leader election protocols. These extractor constructions achieve parameters that go beyond standard resilient functions [AL, Combinatorica’93].
@@ 1 Introduction
With this, we are ready to formally define deterministic condensers:
@@
When $\mathcal{X}$ is the family of all distributions, it is folklore that no non-trivial condensing is possible. [^1] Assuming $m\leq n$ (wlog this holds since $\left\lvert\mathsf{Cond}(\{0,1\}^{n})\right\rvert\leq 2^{n}$), $m-k_{out}\geq(n-k_{in})-\log(1/(1-\varepsilon))$ and hence the output entropy rate cannot be more than the input entropy rate without incurring extremely large error ($>0.999$). So, we additionally assume that $\mathcal{X}$ is a structured family of sources. [^2] A different route, that has been widely studied, is to assume access to a short independent seed. In this work, we will limit ourselves to the deterministic or seedless setting. Since extractors are the highest quality condensers, a significant amount of work has focused on constructing extractors for interesting family of sources, such as: sources generated by small circuits, two independent sources, algebraically generated sources, sources generated by small space sources, and many more [tv00, chattopadhyay_explicit_2019, dgw09extractpolynomials, kamp_deterministic_2007].
@@
In this work, we focus on one natural family of sources where it is known that extraction is impossible (for many interesting parameter regimes). The family we consider are known as online non-oblivious symbol fixing sources (oNOSF sources ). [^3] These sources are in contrast to non-oblivious symbol fixing (NOSF) sources where bad blocks can be arbitrary functions of all the good blocks. NOSF sources were introduced in [chor_bit_1985] with applications in leakage-resilient cryptography, and have been well-studied. Formally:
@@
Our results at a glance
We almost fully resolve the question of the existence of condensers for oNOSF sources by showing that good condensers, meeting the impossibility results of [CGR_seedless_condensers], exist when the number of blocks $\ell$ is growing and $n$ is a large constant. We also construct explicit condensers for oNOSF sources matching the results of [CGR_seedless_condensers] and obtain an improved construction for transforming low-entropy oNOSF sources into uniform oNOSF sources . Moreover, we find new applications of our results on condensers to collective coin flipping and collective sampling, and use these connections to provide simple protocols for these problems. In the context of extractors for oNOSF sources , we introduce the new, natural notion of online influence for Boolean functions and show extraction lower bounds for oNOSF sources by establishing tight bounds on the total online influence of functions. Lastly, using our novel connections to leader election protocols, we construct explicit extractors for uniform oNOSF sources by explicitly constructing the required leader election protocols, the results of which are summarized in Tables 1 and 2.
Organization
The remainder of our introduction is structured as follows. We give an overview of previous work in Section 1.1 before presenting our main existential and explicit condenser results in Section 1.2. In Section 1.3, we present our results on the limits of extraction from oNOSF sources . In Section 1.4, we show how our results on condensers have implications for collective coin flipping and sampling protocols.
@@ 1.1 Previous Work
Extractors
The study of extractors for oNOSF sources was initiated by [aggarwal_how_2020]. [^4] In [aggarwal_how_2020], these sources were called SHELA (Somewhere Honest Entropic Look Ahead) sources. Their results include the following:
  • It is impossible to extract from uniform oNOSF sources when the fraction of good blocks is $0.99$.
@@
Before our work, the best known extractors for oNOSF sources could be obtained by using resilient functions or equivalently, extractors for NOSF sources (non-online version of oNOSF sources ) constructed by [ajtai_influence_1993, chattopadhyay_explicit_2019, meka_explicit_2017, IMV23resilient, IV24resilient] ; these require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$.
Condensers
oNOSF sources were further studied by [CGR_seedless_condensers], where they obtained the following results regarding condensers:
@@
We also mention a related family of sources, namely adversarial Chor-Goldreich sources. Uniform oNOSF sources can be seen as a special case of adversarial Chor-Goldreich sources where the good blocks are uniform. Constructing condensers where the output entropy rate is $g/\ell$ for adversarial Chor-Goldreich sources is already a challenging task, although such condensers in various parameter regimes have been recently constructed [doron_almost_2023, GLZ_cg_condenser]. The paper of [DMOZ24condenseunpredictability] recently constructed condensers for a related more general model.
@@ 1.2.1 Existential Condensers
For all $\ell,\varepsilon$ where $\ell\geq O(\log(1/\varepsilon))$, and $n=10^{4}$, there exists a condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.51\ell,\ell,n\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq 0.99m$ where $m=\Omega(\ell+\log(1/\varepsilon))$. Furthermore, when $n=\omega(1)$, the output entropy rate becomes $1-o(1)$.
@@ 1.2.2 Explicit Condensers
For all $n,\varepsilon$ and constant $\ell$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(0.5\ell+1,\ell\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(m/\varepsilon))$, where $m=\Omega(n)$.
@@ 1.2.3 Transforming Low-Entropy oNOSF sources to uniform oNOSF sources
For all $\ell,n,k,\varepsilon$ where $n=\operatorname{poly}(\log(\ell)),k=O(\log(\ell/\varepsilon))$, there exists a function $f$ such that $f$ transforms ${\left(0.51\ell,\ell,n,k\right)}$-oNOSF sources into uniform ${\left(0.509\ell,\ell,m\right)}$-oNOSF sources with error $\varepsilon$ where $m=\Omega(k)$.
@@
1.3 Limits on Extraction from oNOSF Sources
Next we discuss our results on the limits of extraction from oNOSF sources . Our lower bounds are based on a new notion of influence of functions, namely online influence, that we introduce and analyze. Our upper bound results (explicit extractors) are based on a novel connection to leader election and coin-flipping protocols; to instantiate this connection and give explicit extractors, we construct the necessary protocols.
Extraction Lower bounds via Online Influence
For simplicity, let’s focus on the case of $n=1$, which leads to interesting new questions about Boolean functions. We refer to such uniform ${\left(g,\ell,1\right)}$-oNOSF sources as ${\left(g,\ell\right)}$-oNOBF sources; oNOBF stands for online non-oblivious bit-fixing sources. We ask what is the exact tradeoff between $g$, $\ell$, and $\varepsilon$ for extracting from oNOBF sources . Towards this, we introduce the notion of online influence.
Definition 1.5 (Online influence).
and the total online influence is $\operatorname{\mathbf{oI}}[f]=\sum_{i=1}^{\ell}\operatorname{\mathbf{oI}}_{i}[f].$
We believe this is an interesting new measure and is worth studying in its own right. We refer the reader to Example 7.6 for a couple of interesting examples. For monotone functions (and more generally, unate functions), it is not hard to see that online influence equals the usual notion of influence (see Lemma 7.4 for a proof). Thus, to find interesting properties of online influence (compared to standard influence, Definition 7.1), one must look at non-monotone (in fact, non-unate) Boolean functions.
The following natural question arises towards our goal of proving extractor lower bounds: for a function $f$, what is the maximum online influence out of all $n$ bits? For the usual notion of influence, this question was resolved by the well-known theorem of [kahn_influence_1988], who showed there always exists a bit with influence at least $\operatorname*{Var}(f)\cdot\Omega\left(\frac{\log\ell}{\ell}\right)$.
We show that surprisingly, there exists a balanced function, namely the address function, where every bit has online influence at most $O\left(\frac{1}{\ell}\right)$ (see Lemma 7.12 for a proof). This provides a separation between the usual notion of influence and online influence.
We prove a Poincaré style inequality for total online influence, which shows that the above example is tight. For any function $f:\{0,1\}^{n}\rightarrow\{0,1\}$, define the function $e(f)(x)=(-1)^{f(x)}$.
Theorem 4 (Theorem 7.5, restated).
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(e(f))\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{\ell\operatorname*{Var}(e(f))}$.
Using the above result, we obtain the following extraction lower bound.
Theorem 1.6 (Informal version of Corollary 7.21).
For $\varepsilon<0.01$, there do not exist extractors for ${\left(0.97\ell,\ell\right)}$-oNOBF sources with error at most $\varepsilon$.
Explicit Extractors via Leader Election Protocols
We now move on to presenting our explicit constructions of extractors for oNOBF and oNOSF sources. The following are our main results.
Theorem 5 (informal version of Theorem 8.2).
There exists an explicit function $\mathsf{Ext}:\{0,1\}^{\ell}\to\{0,1\}$ such that for any ${\left(g,\ell\right)}$-oNOBF source $\mathbf{X}$ where $g\geq\ell-\ell/C\log(\ell)$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{1}$, where $C$ is a large constant.
Theorem 6 (informal version of Theorem 8.3).
There exists an explicit function $\mathsf{Ext}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{n}$ such that for any ${\left(g,\ell,n\right)}$-oNOSF source $\mathbf{X}$ where $g\geq\ell-\ell/C\log^{*}(\ell)$ and $n\geq\log(\ell)$, we have $\mathsf{Ext}(\mathbf{X})\approx_{\varepsilon=1/100}\mathbf{U}_{n}$, where $C$ is a large constant.
It is instructive to contrast our results with the non-online setting (where adversarial bits may depend on any good bit). The non-online versions of oNOSF sources and oNOBF sources are called NOSF sources and NOBF sources . For both these sources, the current best extractors require $g\geq\ell-\frac{\ell}{(\log\ell)^{2}}$, which is much more than what our extractors for 5 and 6 require.
We contrast the results for both settings in Tables 1 and 2. In these tables, we are providing known upper and lower bounds on the value of $b(\ell)$, defined as the maximum number of bad symbols for which extraction is still possible with a small constant error — so lower bounds correspond to best known constructions of such functions and upper bounds refers to best known limitation of such functions. We write “$O(\ell)$” to mean “$c\ell$ for some small universal constant $c<1$”.
To interpret our results in terms of (online) influence of coalitions, it will be useful to extend the definition of online influence to subsets of coordinates.
Definition 1.7.
For any function $f:\Sigma^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner. It starts by sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly from $\Sigma$, then picking the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$. Next, the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly from $\Sigma$, and $\mathcal{A}$ sets the value of $x_{i_{2}}$ based on all set variables so for, and so on. Define the advantage of $\mathcal{A}$ to be $\textrm{adv}_{f,B}(\mathcal{A})=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]|$. Then $\operatorname{\mathbf{oI}}_{B}(f)$ is defined to be $\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(\mathcal{A})\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ that control the bits in $B$.
We say a function $f$ is $(b,\varepsilon)$-online-resilient if $\operatorname{\mathbf{oI}}_{B}(f)\leq\varepsilon$ for every set $B\subset[\ell]$ of size at most $b$.
We note that Definition 1.7 is a special case of Definition 1.5, for $\Sigma=\{0,1\}$ and $|B|=1$.
In Section 7.4, we note that online-resilient functions are equivalent to extractors for uniform oNOSF source sources (with one bit output). Thus, our explicit extractor results immediately imply explicit online-resilient functions.
Source Lower bound Upper bound
NOBF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [ajtai_influence_1993] $O\left(\frac{\ell}{\log\ell}\right)$, [kahn_influence_1988]
NOSF $\Omega\left(\frac{\ell}{\log^{2}\ell}\right)$, [ajtai_influence_1993] $O(\ell)$, [bourgain_influence_1992]
Table 1: $b(\ell)$ bounds in the non-online setting.
Source Lower bound Upper bound
oNOBF $\Omega\left(\frac{\ell}{\log\ell}\right)$, [Theorem 8.2] $O(\ell)$, Corollary 7.21 or [aggarwal_how_2020]
oNOSF $\Omega\left(\frac{\ell}{\log^{*}\ell}\right)$, [Theorem 8.3]55footnotemark: 5 $O(\ell)$, [aggarwal_how_2020]
Table 2: $b(\ell)$ bounds in the online setting.
22footnotetext: Recall that this lower bound is for ${\left(g,\ell,n\right)}$-oNOSF sources with $n\geq\log(\ell)$.
Our main technique is a generic way to transform leader election and coin flipping protocols (formally defined in Section 3.3) into extractors for oNOBF and oNOSF sources. This is proved in Lemma 8.1; the general idea of constructing an extractor is to simulate an appropriate leader election protocol with the source at hand (oNOBF or oNOSF), and output according to the chosen leader. To instantiate this transformation, we revisit previous leader election protocols in Section 9. Our leader election protocols provide a slightly stronger than usual guarantee: a good player is elected as the leader with probability close to $1$ (see Lemma 9.1 and Lemma 9.5). This contrasts with the usual guarantee in leader election protocols, where a good leader is chosen with only a non-trivial (constant) probability.
@@ 1.4 Application to Collective Coin Flipping and Collective Sampling
We now discuss applications of our results on condensers for oNOSF sources to fault-tolerant distributed computing. Condensing from oNOSF sources can be viewed as a special case of coin flipping and collective sampling protocols in the full information model that arise in fault-tolerant distributed computing.
@@ 1.4.1 Background
One way to allow for more corruptions (almost linear) among players is to consider “protocols” that allow more rounds of communication. In particular, a protocol can be thought of as a tree where each vertex represents a “round” where in every round the following happens: all good players sends their bits, then all bad players send their bits as a function of the bits of the good players, and they jointly compute a function of these bits. Depending on the outcome of the function, everyone branches on one branch in this tree. Furthermore, every leaf is labeled with final outcomes (say 00 or $1$) and, once a leaf is reached, that is the outcome that everybody agrees on. [ggl98] initiated the study of protocols where the outcomes are from a larger range and where the bad players are trying to minimize the largest probability of any outcome. They called this problem “collective sampling.” For a formal definition, see Section 3.3.
@@ 1.4.2 Known Results
[ben-or_collective_1989] showed that for protocols with outcomes $\{0,1\}$, $b$ bad players can always ensure that some outcome occurs with probability at least $\frac{1}{2}+\frac{b}{2\ell}$. [alonnoar93collective] first constructed a protocol that can handle a linear number of corruptions. Follow-up works tried to reduce the number of rounds in this protocol where, in some settings, players were allowed to send more than one bit per round [russellzuckerman01, feige_noncryptographic_1999].
[ggl98] showed that for all collective sampling protocols and all outcomes, there exists a way for $b$ bad players to coordinate and ensure that an outcome that happens without corruption with probability $p$, now happens with probability $p^{1-(b/n)}\geq p\left(1+\frac{b}{n}\log(1/p)\right)$. Nearly matching collective sampling protocols were constructed by [ggl98, SV08selection, GVZ06selection]. For an overview of further results and bounds, see [dodis2006fault].